if P (2π/3) = ( -1/2 , /3/2) then cos Ø =?



1
Expert's answer
2022-01-16T15:05:23-0500

Solution:

"P (2π/3) = ( -1/2 , /3/2) 

\\\thereforeØ = 2π/3

\\So, \cos Ø = \cos (2π/3)

\\= \cos (π-\dfrac π3)
\\=-\cos (\dfrac π3)
\\=-\dfrac12"